Use the calculator provided to solve the following problems.

  • Consider a t distribution with 15 degrees of freedom. Compute P(t≤−1.29). Round your answer to at least three decimal places.
  • Consider a t distribution with 21 degrees of freedom. Find the value of c such that =P(-c<t<c)=0.90 Round your answer to at least three decimal places.
Answer #1

Solution:-

Given that,

Degrees of freedom = 15

P( t < -1.29)

= 0.1083

Degrees of freedom = 21

P( -c < t < c) = 0.90

= P(t < c) - P(Z <-z ) = 0.90

= 2P(t < c) - 1 = 0.90

= 2P(t < c) = 1 + 0.90

= P(t < c) = 1.9 / 2

= P(t < c) = 0.95

= P(t < 1.721) = 0.95

= c ± 1.721